OpenStudy (anonymous):

Explain how you would solve the equation -7 + x = 12 using the properties of equality

OpenStudy (accessdenied):

Do you know the properties of equality?

OpenStudy (anonymous):

Yes

OpenStudy (accessdenied):

Well, since our goal is to get the 'x' by itself, we are able to use addition to move the other term, '-7', to the other side while cancelling the one next to the x. -7 + x = 12 -7 + x + 7 = 12 +7 Addition prop. of eq. x = 19 simplification